OpenStudy (anonymous):

what is the value of X when cos(x)=1\2

OpenStudy (anonymous):

\[\LARGE x=\cos^{-1} (\frac{1}{2})\]

OpenStudy (anonymous):

\(\Large \cos^{-1} (\frac{1}{2})\) means the angle at which cosine gives 1/2 Do you know any such angle?

OpenStudy (anonymous):

yes x is equal at 60 degrees

OpenStudy (anonymous):

yepp :)